# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ 4.4 Midplane accretion structure in the disc
The accretion flow at the midplane is controlled by the stress profiles in the disc. Fig. 10 shows radial profiles of $\alpha$ and $\alpha_{\mathcal{M}}$ at the midplane. In the inner MRI active zone, $\alpha\sim 3\times{10}^{-3}$ with $\alpha_{\mathcal{M}}/\alpha_{\mathcal{R}}\sim 3$, consistent with typical MRI-saturated dynamo behaviour. Furthermore, $\alpha_{\mathcal{M}}$ at the midplane is $\approx 40$ per cent lower compared to the volume-averaged value over the MRI active region shown in Fig. 4. This is the result of the inhomogeneous vertical magnetic stress structure in the active zone, shown in Fig. 8.
@@
Meanwhile, in the inner dead zone, the total midplane $\alpha$ is $\sim 5\times 10^{-4}$ which is nearly a full order of magnitude lower than in the active zone. This stress is driven almost entirely by the Reynolds component [^2] Section 6.2 extends this analysis to the vertical Maxwell stress profile in the inner dead zone, driven by large-scale magnetic field structures., attributed to large-scale density waves emanating from vortex structures (Section 5.2; Chametla et al. 2024), fluctuations at the dead–active zone interface (Faure et al., 2014) and the MRI active zone (Heinemann & Papaloizou, 2009). These waves decay with radius due to dissipation by weak shocks (e.g. Faure et al., 2014), leading to a decrease in the associated Reynolds $\alpha$.
Figure 10: Radial profiles of $\alpha$ (black), the Maxwell $\alpha$ (red) and the accretion flow, quantified by $\langle\rho u_{r}\rangle_{\phi,t}$ (blue). The quantities are evaluated at the midplane and have been averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$. The grey shaded region highlights that the midplane flow is outward at the outer edge of the MRI active zone.
@@
The midplane accretion flow is quantified with $\langle\rho u_{r}\rangle_{\phi,t}$, and its radial profile is shown in Fig. 10. In the inner MRI active region ($R\lesssim 0.75$ au) and throughout the dead zone, the time-averaged flow is inward. However, at the outer edge of the MRI dead region $(0.75\,\text{au}\lesssim R\lesssim 0.95\,\text{au})$ there is sustained outward flow at the midplane. This configuration is present in all four dead–active zone interface simulations and persists throughout their evolution. We attribute this outward flux to the radial gradient of the torque distribution, noting that the direction of radial mass flux is set by the sign of $-\partial_{R}(R^{2}\alpha P)$, in the absence of strong vertical stresses (see Lesur, 2021a).
@@ 5.1 Pressure Bump
A coherent, local axisymmetric density maximum forms just outside the dead–active zone interface in ZNF-FID, developing rapidly within a few local orbits. To test the any dependence of this rapid evolution on the initial conditions, we perform the ZNF-MRI-IC simulation, in which the non-ideal MHD profiles are imposed on top of a fully MRI turbulent disc. The pressure bump still forms in a comparable manner after these non-ideal terms are introduced, indicating that this outcome is not sensitive to the specific initial setup.
@@
Radial profiles of the surface density $\Sigma(R)$ during the early stages of evolution are shown in the top panel of Fig. 11. These profiles demonstrate that the local maximum forms due to a density enhancement at the inner edge of the dead zone, rather than solely from the inward draining of material in the MRI active zone.
@@
Figure 11: Radial profiles of the surface density $\Sigma$ over the initial six orbits ($t_{\text{in}}=48$) at the dead–active zone interface (dashed grey line). Note the similarity in evolution between the three-dimensional MHD simulation ZNF-MRI-IC (top) and the one-dimensional viscous HD model (bottom), and the density enhancement around the dead–active zone interface. In the top panel, $t_{\text{DZI}}=0$ references the time that the non-ideal MHD model is placed on top of a fully MRI turbulent disc, and $\Sigma$ is smoothed over a small temporal average ($t_{\text{in}}=0.5$).
@@
To investigate this process, we employed the one-dimensional viscous disc diffusion formalism introduced in Section 3.4.4. The setup assumed an initial Keplerian configuration, as given in equation (26), and an initial surface density scaling of $\Sigma\propto R^{-1/2}$, consistent with the initial condition (10). Fixed radial boundary conditions were applied to mimic the buffer zones used in the three-dimensional simulations. The formalism was closed by setting a vertically integrated radial profile of $\alpha_{\mathcal{M}}$, consistent with the midplane profile shown in Fig. 10, thereby defining a steady profile for the kinematic viscosity $\bar{\nu}$. We then solved the system on a radial grid comparable to the that used in ZNF-MRI-IC; the resulting evolution of $\Sigma$ is shown in the bottom panel of Fig. 11.
@@
The one-dimensional model closely reproduces the evolution observed in ZNF-MRI-IC over the first six local orbits, confirming that the density enhancement at the dead–active zone interface results from a radial viscous process. To further quantify this behaviour, Fig. 12 shows the initial radial profiles of the average viscous torque $\mathcal{G}$ and the corresponding mass flux $\mathcal{F}$, as defined in equations (27) and (28), using the steady-state $\bar{\nu}$ profile. The results show that the density enhancement originates from a negative radial gradient in viscous torque ($\partial_{R}\mathcal{G}<0$), which drives an outward mass flux from the outer edge of the active zone into the dead zone.
@@
Figure 12: Radial profiles of the initial average viscous torque $\mathcal{G}$ (blue) and mass flux $\mathcal{F}$ (red) for the one-dimensional viscous disc diffusion model, outlined in Section 5.1. The grey shaded region denotes where the mass flux is outward $(\mathcal{F}>0)$, and it explains the formation of local mass enhancement just outside the dead–active zone interface (dashed grey line) shown in Fig. 11.
@@
In essence, this outward drift resembles the outward expansion of the outer edge of a finite-sized viscous accretion disc. The material near the dead–active zone interface receives angular momentum from material interior to it, but due to the lower $\alpha$ in the dead zone, lacks an efficient mechanism to transport that momentum further outward. Consequently, this material itself must move outward. In this way, the dead–active zone interface effectively acts as a barrier to outward angular momentum transport. Furthermore, despite being governed by viscous processes, the short radial length scales $\delta R$ involved, mean that the local viscous timescale $\tau_{\text{visc}}$ can approach the dynamical timescale, since $\tau_{\text{visc}}\sim\delta R^{2}/\bar{\nu}\sim\alpha^{-1}(\delta R/H)^{2}\Omega_{K}^{-1}$.
@@
Interestingly, both panels of Fig. 11 reveal a fixed point, where the net radial mass flux is zero such that $\Sigma$ is constant. This coincides with a turning point in the mass flux profile (red line in Fig. 12), highlighting a generic feature of the initial flow structure in this region.
@@
The axisymmetric pressure maximum is sustained throughout the duration of the simulations and remains vertically coherent through $z\approx\pm 3.5H$. This is illustrated (up to $z=\pm 3H$) by the normalised radial pressure profiles $\langle P\rangle_{\phi,t}/\langle P(r=r_{\text{in}})\rangle_{\phi,t}$ along slices of constant $\theta$ shown in Fig. 13. Notably, this coherence persists despite the complex vertical structure of the MRI active zone discussed in Section 4.3.
It appears that the sufficiently strong ambipolar diffusion profile (see Fig. 1) suppresses MRI activity in the surface layers of the dead zone, unlike in the Ohmic-resistive three-dimensional simulations of Flock et al. (2017), thereby enabling the formation of this vertically extended pressure structure. This has important implications for dust dynamics, as it can reduce the radial drift of solids across the vertical extent of the disc at the dead–active zone interface.
Figure 14: Snapshots of the residual vertical vorticity $(\omega_{z}-\omega_{z_{K}})$ at the disc midplane for ZNF-FID. From top to bottom, the dominant azimuthal mode structures are $m=7$, $m=5$, $m=2$ and $m=1$, in agreement with Fig. 17. The anticyclonic vortices encroach upon the MRI active region (the dashed black line denotes the dead–active zone interface) and the eventual $m=1$ vortex possesses a weak and disrupted spatial morphology (bottom panel) during its lifetime.
@@ 5.2.1 Formation
Anticyclonic vortices form at the dead–active zone interface in all four dead–active zone interface simulations. This is illustrated for ZNF-FID in Fig. 14, which shows snapshots of the residual vertical vorticity $(\omega_{z}-\omega_{z_{K}})$ at the midplane, where $\omega_{z_{K}}=\Omega_{K}/2$ is the Keplerian vertical vorticity.
@@
These vortices arise from the saturation of the RWI (Lovelace et al., 1999; Lovelace & Romanova, 2014). The variation of RWI modes in $\phi$ can be captured with an azimuthal Fourier decomposition, characterised by the wavenumber $m$, as defined in equation (17). In our setup, the $m=7$ mode grows fastest, initially forming seven coherent anticyclonic vortices (see top panel of Fig. 14).
@@
To compare with linear theory, we examine the Fourier power time series of the azimuthal velocity deviation $u_{\phi}^{\prime}$, defined in equation (19), evaluated at the vortex centres: $P_{u_{\phi}^{\prime},m}\,(R=1.05\,\mathrm{au},t)$. This is computed across modes with $m<10$, as defined in equation 17. Fig. 15 shows the initial power time series for the $m=7$ mode, which undergoes exponential growth followed by saturation at $t_{\text{in}}\approx 50$, consistent with the residual vertical vorticity plots in Fig. 14. The growth rate of the $m=7$ mode, corresponding to a perturbation $\propto\exp(\sigma t)$, is $\sigma\sim 0.17\,\Omega_{K}(R=1.05\,\text{au})$, in qualitative agreement with (Li et al., 2000). Additional measurements show slower growth rates for smaller and larger $m$. In contrast with Li et al. (2001), we do not find the $m=4$ or $m=5$ modes to be the fastest growing. Instead, the rapid reshaping of the vortensity profile at the interface produces a sharper, narrower minimum, favouring a higher $m=7$ mode. Ono et al. (2016) showed that dominance of the $m=7$ mode places the system near the Rayleigh stability threshold under their assumptions; however, it remains stable throughout our simulation, despite the sharp interface.
@@
Figure 15: Fourier power time series of the azimuthal velocity deviation from Keplerian flow $P_{u_{\phi}^{\prime},m=7}\,(R=1.05\,\text{au},t)$, for the fastest growing Rossby wave mode ($m=7$), evaluated at the vortex centres. The grey shaded region indicates the linear growth phase of the RWI, during which the approximate growth rate is $\sigma\sim 0.17\Omega_{K}$ for the $m=7$ mode.
@@
Whilst there is no general analytic criterion for the onset of the RWI, it is the presence of a vortensity minimum, and not a pressure maximum that is critical. We confirm this in Fig. 16, which shows radial profiles of the pressure $\langle P\rangle_{\phi}$ and vortensity $\langle\xi\rangle_{\phi}$, defined in equation (18), at $t_{\text{in}}=10$, corresponding to the onset of the RWI’s linear phase (see Fig. 15). A distinct localised vortensity minimum is present just outside the dead–active zone interface, with minimal variation in the pressure structure.
@@
Rapid onset of the RWI is also confirmed in ZNF-MRI-IC, demonstrating that it is not simply a consequence of the initial $\rho$ or $\mathbf{B}$ profiles. Moreover, we find consistency with the empirical criterion proposed by Chang et al. (2023), which predicts that the RWI is triggered when the radial oscillation frequency – the epicyclic frequency for a barotropic flow – is $\lesssim 75$ per cent of the local Keplerian frequency.
@@
Figure 16: Radial profiles of the azimuthally averaged vortensity $\langle\xi\rangle_{\phi}$ (red) and pressure $\langle P\rangle_{\phi}$ (black) at the disc midplane. The profiles are taken at $t_{\text{in}}=10$, which corresponds to the start of the linear growth phase of the RWI (see Fig. 15). Close to the dead–active zone interface (dashed grey line) there is a distinct minimum in the vortensity and no local pressure maximum. The pressure profile is normalised by a factor of $10^{3}$.
@@ 5.2.2 Initial evolution and vortex morphology
Radial profiles of the Fourier power time series $P_{u_{\phi}^{\prime},m}$, for the azimuthal mode numbers $m\in\{1,2,3,4,5\}$, are shown in Fig. 17. Whilst, the $m=7$ mode dominates initially, its contribution is negligible after $t_{\text{in}}\approx 65$ and is omitted for clarity. The power distribution then sequentially evolves through coherent $m=5$ and $m=2$ phases, consistent with the snapshots shown in Fig. 14, before evolving into an $m=1$ structure by $t_{\text{in}}\approx 230$. This corresponds to a vortex merger timescale of approximately $20$ local orbits.
Figure 17: Fourier power time series of the azimuthal velocity deviation from Keplerian flow, $P_{u_{\phi}^{\prime},m}\,(R=1.05,\text{au},t)$, evaluated at the vortex centres for azimuthal mode numbers $m\in\{1,2,3,4,5\}$. Compare the vortex merger process shown here with Fig. 14, and note the subsequent decay of the power budget associated with the weakly coherent $m=1$ structure. The power is normalised by a factor of $10^{6}$.
The $m=1$ anticyclonic vortex possesses a weak and spatially incoherent morphology (see bottom panel of Fig. 14). We attribute this weakened structure to the turbulent dynamics and associated wave excitation within the MRI active zone (e.g. Heinemann & Papaloizou, 2009), into which the nearby vortices encroach.
Additionally, the vortices do not migrate and remain fixed at the localised pressure maximum, in agreement with Paardekooper et al. (2010). This behaviour is different from the migrating vortices seen in more thermodynamically complex simulations by Faure et al. (2015).
Finally, through $m$-mode analysis of the disc midplane, we establish a connection between the $m$-structured density waves in the dead zone and the vortices, as the waves retain the asymmetric structure of the vortices that launch them. These waves contribute to the midplane $\alpha_{\mathcal{R}}$ on the order of $10^{-4}$ in the dead zone shown in Fig. 10, offering a weak mechanism for angular momentum transport. They may also aid in the vertical lofting of dust and intermittently concentrate solids within their crests.

### v3

@@ 4.4 Midplane accretion structure
The accretion flow at the midplane is controlled by the stress profiles in the disc. Fig. 10 shows radial profiles of $\alpha$ and $\alpha_{\mathcal{M}}$ at the midplane. In the inner MRI-active zone, $\alpha\!\sim\!3\times{10}^{-3}$ with $\alpha_{\mathcal{M}}/\alpha_{\mathcal{R}}\!\sim\!3$, consistent with typical MRI-saturated dynamo behaviour. Furthermore, $\alpha_{\mathcal{M}}$ at the midplane is roughly $40$ per cent lower compared to the volume-averaged value over the MRI-active region shown in Fig. 4. This is the result of the inhomogeneous vertical magnetic stress structure in the active zone, shown in Fig. 8.
@@
Meanwhile, in the inner dead zone, the total midplane $\alpha$ is $\sim\!5\times 10^{-4}$, which is nearly an order of magnitude lower than in the active zone. This stress is driven almost entirely by the Reynolds component, attributed to large-scale density waves emanating from vortex structures (see Section 5.2; chametla_turbulent_2024), fluctuations at the dead–active zone interface (faure_thermodynamics_2014), and the MRI-active zone (heinemann_excitation_2009). [^4] Section 6.2 extends this analysis to the vertical Maxwell stress profile in the inner dead zone, driven by large-scale magnetic field structures. These waves decay with radius due to dissipation by weak shocks (e.g. faure_thermodynamics_2014), leading to a decrease in the associated Reynolds $\alpha$.
@@
The midplane accretion flow is quantified with $\langle\rho u_{r}\rangle_{\phi,t}$, and its radial profile is shown in Fig. 10. In the inner MRI-active region ($R\lesssim 0.75$ au) and throughout the dead zone, the time-averaged flow is inward. However, at the outer edge of the MRI-active region $(0.75\,\text{au}\lesssim R\lesssim 0.95\,\text{au})$ there is sustained outward flow at the midplane. This configuration is present in all four dead–active zone interface simulations and persists throughout their evolution. We attribute this outward flux to the radial gradient of the torque distribution, noting that the direction of radial mass flux is set by the sign of $\partial_{R}(R^{2}\alpha P)$, in the absence of strong vertical stresses (see lesur_magnetohydrodynamics_2021).
@@
Figure 10: Radial profiles of $\alpha$ (black), $\alpha_{\mathcal{M}}$ (red), and the accretion flow, quantified by $\langle\rho u_{r}\rangle_{\phi,t}$ (blue). The quantities are evaluated at the midplane and have been averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$. The grey-shaded region highlights that the midplane flow is outward at the outer edge of the MRI-active zone.
@@ 5.1 Pressure bump
A coherent, local axisymmetric density maximum forms just outside the dead–active zone interface in ZNF-FID, developing rapidly within a few local orbits. To test any dependence of this rapid evolution on the initial conditions, we perform the ZNF-MRI-IC simulation, in which the non-ideal MHD profiles are imposed on top of a fully MRI-turbulent disc. The pressure bump still forms in a comparable manner after these non-ideal terms are introduced, indicating that this outcome is not sensitive to the specific initial setup.
@@
Radial profiles of the surface density, $\Sigma(R)$, during the early stages of evolution are shown in the top panel of Fig. 11. These profiles demonstrate that the local maximum forms due to a density enhancement at the inner edge of the dead zone, rather than solely from the inward draining of material in the MRI-active zone.
@@
Figure 11: Radial profiles of the surface density, $\Sigma$, over the initial six orbits ($t_{\text{in}}=48$) at the dead–active zone interface (dashed grey line). Note the similarity in evolution between the three-dimensional MHD simulation ZNF-MRI-IC (top) and the one-dimensional viscous HD model (bottom), and the density enhancement around the dead–active zone interface. In the top panel, $t_{\text{DZI}}=0$ references the time that the non-ideal MHD model is placed on top of a fully MRI-turbulent disc, and $\Sigma$ is smoothed over a small temporal average of $t_{\text{in}}=0.5$.
@@
To investigate this process, we used the one-dimensional viscous-disc diffusion formalism introduced in Section 3.4.4. The setup assumed an initial Keplerian configuration, as given in equation (25), and an initial surface density of $\Sigma\propto R^{-1/2}$, consistent with the initial condition (10). Fixed radial boundary conditions were applied to mimic the buffer zones used in the three-dimensional simulations. The formalism was closed by setting a vertically integrated radial profile of $\alpha_{\mathcal{M}}$, consistent with the midplane profile shown in Fig. 10, thereby defining a steady profile for the kinematic viscosity, $\bar{\nu}$. We then solved the system on a radial grid comparable to that used in ZNF-MRI-IC; the resulting evolution of $\Sigma$ is shown in the bottom panel of Fig. 11.
@@
The one-dimensional model closely reproduces the evolution observed in ZNF-MRI-IC over the first six local orbits, confirming that the density enhancement at the dead–active zone interface results from a radial viscous process. To further quantify this behaviour, Fig. 12 shows the initial radial profiles of the average viscous torque, $\mathcal{G}$, and the corresponding mass flux, $\mathcal{F}$, as defined in equations (26) and (27), using the steady-state $\bar{\nu}$ profile. The results show that the density enhancement originates from a negative radial gradient in viscous torque ($\partial_{R}\mathcal{G}<0$), which drives an outward-directed mass flux from the outer edge of the active zone into the dead zone.
@@
Figure 12: Radial profiles of the initial average viscous torque, $\mathcal{G}$ (blue), and mass flux, $\mathcal{F}$ (red), for the one-dimensional viscous-disc diffusion model, outlined in Section 5.1. The grey-shaded region denotes where the mass flux is outward $(\mathcal{F}>0)$, explaining the formation of a local mass enhancement just outside the dead–active zone interface (dashed grey line) shown in Fig. 11.
@@
In essence, this outward drift resembles the expansion of the outer edge of a finite-sized viscous accretion disc. The material near the dead–active zone interface receives angular momentum from material interior to it, but due to the lower $\alpha$ in the dead zone, lacks an efficient mechanism to transport that momentum further outward. Consequently, this material itself must move outward. In this way, the dead–active zone interface effectively acts as a barrier to outward angular momentum transport. Furthermore, despite being governed by viscous processes, the short radial length scales, $\delta R$, involved, mean that the local viscous timescale, $\tau_{\text{visc}}$, can approach the dynamical timescale, since $\tau_{\text{visc}}\!\sim\!\delta R^{2}/\bar{\nu}\!\sim\!\alpha^{-1}(\delta R/H)^{2}\Omega_{\text{K}}^{-1}$.
@@
Interestingly, both panels of Fig. 11 reveal a fixed point where the net radial mass flux is zero, so that $\Sigma$ is constant. This coincides with a turning point in the mass flux profile (red line in Fig. 12), highlighting a generic feature of the initial flow structure in this region.
@@
The axisymmetric pressure maximum is sustained throughout the duration of the simulations and remains vertically coherent through $z\approx\pm 3.5H$. This is illustrated (up to $z=\pm 3H$) by the normalised radial pressure profiles, $\langle P\rangle_{\phi,t}/\langle P(r=0.25\,\text{au})\rangle_{\phi,t}$, along the slices of constant $\theta$ shown in Fig. 13. Notably, this coherence persists despite the complex vertical structure of the MRI-active zone discussed in Section 4.3.
It appears that the sufficiently strong ambipolar diffusion profile (see Fig. 1) suppresses MRI activity in the surface layers of the dead zone, unlike in the Ohmic-resistive three-dimensional simulations of flock_3d_2017, thereby enabling the formation of this vertically extended pressure structure. This has important implications for dust dynamics, as it can reduce the radial drift of solids across the vertical extent of the disc at the dead–active zone interface.
Figure 14: Snapshots of the residual vertical vorticity $(\omega_{z}-\omega_{z_{\text{K}}})$ at the disc midplane for ZNF-FID. From top to bottom, the dominant azimuthal mode structures are $m=7$, $m=5$, $m=2$ and $m=1$, in agreement with Fig. 17. The anticyclonic vortices encroach upon the MRI-active region (the dashed black line denotes the dead–active zone interface) and the eventual $m=1$ vortex possesses a weak and disrupted spatial morphology (bottom panel) during its lifetime.
@@ 5.2.1 Formation
Anticyclonic vortices form at the dead–active zone interface in all four dead–active zone interface simulations. This is illustrated for ZNF-FID in Fig. 14, which shows snapshots of the residual vertical vorticity $(\omega_{z}-\omega_{z_{\text{K}}})$ at the midplane, where $\omega_{z_{\text{K}}}=\Omega_{\text{K}}/2$ is the Keplerian vertical vorticity.
@@
These vortices arise from the saturation of the RWI (lovelace_rossby_1999; lovelace_rossby_2014). The variation of RWI modes in $\phi$ can be captured with an azimuthal Fourier-mode decomposition, characterised by the wave number $m$, as defined in equation (17). In our setup, the $m=7$ mode grows fastest, initially forming seven coherent anticyclonic vortices (see top panel of Fig. 14).
@@
To compare with linear theory, we examine the Fourier-power time series of the azimuthal velocity deviation, $u_{\phi}^{\prime}$, defined in Section 3.4.2, evaluated at the vortex centres: $P_{u_{\phi}^{\prime},m}\,(R=1.05\,\mathrm{au},t)$. This is computed across modes with m<10m\!<\!10, as defined in equation (17). Fig. 15 shows the initial power time series for the $m=7$ mode, which undergoes exponential growth followed by saturation at $t_{\text{in}}\!\sim\!50$, consistent with the residual vertical vorticity plots in Fig. 14. The growth rate of the $m=7$ mode, corresponding to a perturbation $\propto\exp(\sigma t)$, is $\sigma\!\sim\!0.17\,\Omega_{\text{K}}(R=1.05\,\text{au})$, which is roughly consistent with li_rossby_2000. Additional measurements show slower growth rates for smaller and larger $m$. In contrast with li_rossby_2001, we do not find the $m=4$ or $m=5$ modes to be the fastest growing. Instead, the rapid reshaping of the vortensity profile at the interface produces a sharper, narrower minimum, favouring a higher $m=7$ mode. ono_parametric_2016 showed that dominance of the $m=7$ mode places the system near the Rayleigh stability threshold under their assumptions; however, it remains stable throughout our simulation, despite the sharp interface.
@@
Figure 15: Fourier-power time series of the azimuthal velocity deviation from Keplerian flow $P_{u_{\phi}^{\prime},m=7}\,(R=1.05\,\text{au},t)$, for the fastest growing Rossby wave mode ($m=7$), evaluated at the vortex centres. The grey-shaded region indicates the linear growth phase of the RWI, with an approximate growth rate of $\sigma\sim 0.17\Omega_{\text{K}}$.
@@
Whilst there is no general analytic criterion for the onset of the RWI, it is the presence of a vortensity minimum, and not a pressure maximum that is critical. We confirm this in Fig. 16, which shows radial profiles of the pressure, $\langle P\rangle_{\phi}$, and vortensity, $\langle\xi\rangle_{\phi}$, defined in equation (18), at $t_{\text{in}}=10$, corresponding to the onset of the RWI’s linear phase (see Fig. 15). A distinct localised vortensity minimum is present just outside the dead–active zone interface, with minimal variation in the pressure structure.
@@
The rapid onset of the RWI is also confirmed in ZNF-MRI-IC, demonstrating that it is not simply a consequence of the initial $\rho$ or $\mathbf{B}$ profiles. Moreover, we find consistency with the criterion proposed by chang_origin_2023, which predicts that the RWI is triggered when the radial oscillation frequency – the epicyclic frequency for a barotropic flow – is $\lesssim 75$ per cent of the local Keplerian frequency.
@@
Figure 16: Radial profiles of the azimuthally averaged vortensity, $\langle\xi\rangle_{\phi}$ (red), and pressure, $\langle P\rangle_{\phi}$ (black), at the midplane. The profiles are taken at $t_{\text{in}}=10$, which corresponds to the start of the linear growth phase of the RWI (see Fig. 15). Close to the dead–active zone interface (dashed grey line) there is a distinct minimum in the vortensity and no local pressure maximum. The pressure profile is normalised by a factor of $10^{3}$.
@@ 5.2.2 Initial evolution and vortex morphology
Radial profiles of the Fourier-power time series, $P_{u_{\phi}^{\prime},m}$, for the azimuthal mode numbers $m\in\{1,2,3,4,5\}$, are shown in Fig. 17. Whilst, the $m=7$ mode dominates initially, its contribution is negligible after $t_{\text{in}}\!\sim\!65$ and is omitted for clarity. The power distribution then sequentially evolves through coherent $m=5$ and $m=2$ phases, consistent with the snapshots shown in Fig. 14, before evolving into an $m=1$ structure by $t_{\text{in}}\!\sim\!230$. This corresponds to a vortex-merger timescale of approximately $20$ local orbits.
Figure 17: Fourier-power time series of the azimuthal velocity deviation from Keplerian flow, $P_{u_{\phi}^{\prime},m}\,(R=1.05,\text{au},t)$, evaluated at the vortex centres for azimuthal mode numbers $m\in\{1,2,3,4,5\}$. Compare the vortex succession process shown here with Fig. 14, and note the subsequent decay of the power budget associated with the weakly coherent $m=1$ structure. The power is normalised by a factor of $10^{6}$.
The $m=1$ anticyclonic vortex possesses a weak and spatially incoherent morphology (see bottom panel of Fig. 14). We attribute this weakened structure to the turbulent dynamics and associated wave excitation within the MRI-active zone (e.g. heinemann_excitation_2009), into which the nearby vortices encroach.
Additionally, the vortices do not migrate and remain fixed at the localised pressure maximum, in agreement with paardekooper_vortex_2010. This behaviour is different from the migrating vortices seen in more thermodynamically complex simulations by faure_vortex_2015.
Finally, through $m$-mode analysis of the disc midplane, we establish a connection between the spiral density waves in the dead zone and the vortices, as at any instant, the waves exhibit the same $m$ as the vortices that launch them. These waves contribute to the midplane $\alpha_{\mathcal{R}}$ on the order of $10^{-4}$ in the dead zone shown in Fig. 10, offering a weak mechanism for angular momentum transport. They may also aid in the vertical lofting of dust and intermittently concentrate solids within their crests.